# MeGA: Merging Multiple Independently Trained Neural Networks Based on Genetic Algorithm

### v2

@@ 4.2 CIFAR-10 Results
Table 1: Performance Comparison of Individual and MeGA-Merged CNN Models based on MeGA with CIFAR-10 Dataset
Dataset Model (# of Params.) Baseline Baseline Acc. Final Test Acc. (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (56) 0.920 0.945 Model 2 (57) 0.942 SVHN [12] ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 (10) Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 Model 2 (57) 0.781 ResNet 56 (0.86M) [12] Model 1 (56) 0.830 0.862 Model 2 (57) 0.860 Caltech-101 [12] ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 (10) Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 Model 2 (57) 0.781 ResNet 56 (0.86M) [12] Model 1 (56) 0.830 0.862 Model 2 (57) 0.860 SVHN [12] ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 (10) Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 Model 2 (57) 0.781

### v3

@@ 4.2 CIFAR-10 Results
Table 1: Performance Comparison of Individual and Merged CNN Models based on MeGA with CIFAR-10 Dataset
Datasets Model (# of Params.) Baseline. Baseline Test Acc. Final Test Acc. # of classes (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (56) 0.809 0.822 Model 2 (57) 0.793 ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 CIFAR-10 [18] Model 2 (57) 0.781 (10) Xception (22.96M) [12] Model 1 (56) 0.716 0.754 Model 2 (57) 0.730 DenseNet 121 (7.04M) [12] Model 1 (56) 0.719 0.742 Model 2 (57) 0.728 DenseNet 169 (12.65M) [12] Model 1 (56) 0.735 0.753 Model 2 (57) 0.731